OpenStudy (anonymous):

Which of the following equations best represents the graph of a function f(m)? f(y) = m f(x) = m y = f(m) x = f(m)

OpenStudy (anonymous):

@cj49 @gorv @shinalcantara

OpenStudy (gorv):

but it depends in what context you are asking

OpenStudy (gorv):

generally we put value of x

OpenStudy (gorv):

and calculate y from that

OpenStudy (gorv):

sorry y=f(m) it will be

OpenStudy (cj49):

it wud be y=f(m)

OpenStudy (gorv):

as we know y=f(x)

OpenStudy (gorv):

at x=m y=f(m)

OpenStudy (anonymous):

Thanks alot! can you help me with a couple more?

OpenStudy (gorv):

yeah sure

OpenStudy (anonymous):

ill open it up in another post so i can give more medals